OpenStudy (anonymous):

If f(x)=5x+8 and get g(x)=4x-1, find f(g(x))

OpenStudy (amistre64):

insert g(x) into the spot of "x" within f(x)

OpenStudy (amistre64):

f(g(x))=5(g(x))+8 and g(x)=4x-1 sooo.... f(g(x))=5(4x-1)+8 and simplify as wanted

OpenStudy (anonymous):

that's what I though. Thanks :)

OpenStudy (anonymous):

sooo.... f(g(x))= 20x+3

OpenStudy (anonymous):

now if I'm finding that for g(f(3)) what do I do?

OpenStudy (anonymous):

basically same thing. plug in 3 for value of x once u find g(f(x)) 4(5(3)+8)-1, 4(15+8)-1, 4(23)-1, 92-1=91